OpenStudy (anonymous):

in a certain oscillatory system, the amplitude of motion is 5m and the time period is 4s. the time taken by the particle for passing between points which are at distances of 4m and 2m from the centre and on the same side of it will be?

OpenStudy (anonymous):

hmm.

OpenStudy (anonymous):

pls help me find an answer to this one... i had spentt a lot of time in it...asked to many..

OpenStudy (anonymous):

What do your notes say?

OpenStudy (anonymous):

Can you not find anything in your notes, or can you not understand the passage?

OpenStudy (anonymous):

i got this question from "NCERT AT YOUR FINGERTIPS : PHYSICS" by mtg. the problem is for this question , there is no explanation..just the answer which is 0.33 sec. and the options were 0.31,0.32,0.33,0.34 ....

OpenStudy (anonymous):

ok use equation \[x= asin \omega t\] you can calculate \[\omega = 2 \pi / T\] now calculate t1 and t2 as\[x _{1}= asin \omega t _{1} and x _{2}= asin \omega t _{2}\] calculate t1 and t2 subtract you will get it.

OpenStudy (anonymous):

ya..thank you for your reply...i also thought this way but the problem is that we can't know the value of \[ \sin^{-1} \frac{ 4 }{ 5 } \] and \[\sin^{-1} \frac{ 2 }{ 5 }\] directly without using log book. for entrance exam log books aren't allowed...so i thought there could be some other method to do this...

OpenStudy (anonymous):

may be but i don't know.